OpenStudy (anonymous):

Algebra Help Plz Will Fan & Medal ;) 1. Simplify the sum or difference. sqrt 5 + 6 sqrt 5 A) 6 sqrt 5 B) 30 C) 6 sqrt 10 D) 7 sqrt 5 2. Simplify the sum or difference. 4 sqrt 2 - 7 sqrt 2 A) negative sqrt 6 B) -3 sqrt 2 C) -11 sqrt 2 D) 11 sqrt 2 3. Simplify the sum or difference. 3 sqrt 7 - sqrt 63 A) -2 sqrt 7 B) -6 sqrt 7 C) -4 sqrt 3 D) 0

OpenStudy (anonymous):

@jacobciezki

OpenStudy (johnweldon1993):

Whatcha got for answers that you've tried? :)

rebeccaxhawaii (rebeccaxhawaii):

caught in your tracks john HAH math cheater

OpenStudy (johnweldon1993):

XDD hey when you need help again I got you @rebeccaxhawaii

rebeccaxhawaii (rebeccaxhawaii):

ikik im jk thanks

OpenStudy (johnweldon1993):

Okay sorry about that @Answers101 haha...didnt mean to spam your question But yeah, I'm assuming you have tried these, what do you have so far?

OpenStudy (anonymous):

@johnweldon1993 1. A 2. C 3. A

OpenStudy (johnweldon1993):

2 is good :) 1...not quite...if we have \(\large \sqrt{5} + \sqrt{5}\) what does that equal? \(\large 2\sqrt{5}\) right? So applying that here...we have \(\large \sqrt{5} + 6\sqrt{5} = 7\sqrt{5}\) right?

OpenStudy (anonymous):

It's really confusing to me :(

OpenStudy (johnweldon1993):

Okay...well hmm lets try and think about it a little bit before hand Lets write some general rules \[\large 1\sqrt{a} + 1\sqrt{a} = 2\sqrt{a}\] \[\large 2\sqrt{a} + 3\sqrt{a} = 5\sqrt{a}\] So what do you notice? When the number under the square root sign is the same...we leave it alone and just add the numbers outside together \[\large 7\sqrt{a} + 10\sqrt{a} = 17\sqrt{a}\] idk if that helps at all...

OpenStudy (kobi):

what number

OpenStudy (johnweldon1993):

You actually already did it for 2) \[\large 4\sqrt{2} - 7\sqrt{2}\] since they both have \(\large \sqrt{2}\) we just leave that alone...and worry about the numbers on the outside \(\large 4\sqrt{2} - 7\sqrt{2} \rightarrow (4 - 7)\sqrt{2}\) which = \(\large -11\sqrt{2}\)